Computing Present Values of Single Amounts and Annuities
Refer to Tables 1 and 2 in Appendix A near the end of the book to compute the present value for each of the following amounts.

Round answers to the nearest dollar.

a. $130,000 received 10 years hence if the annual interest rate is:

10% compounded annually
10% compounded semi-annually


d. $260,000 received 10 years hence discounted at 10% per year compounded annually.  
Answer #1

Solution a1:

Present value = $130,000 * PV factor at 10% for 10th period

= $130,000 * 0.38554 = $50,120

Solution a2:

Semiannual interest rate = 10%/2 = 5%

Nos of semiannual period = 10*2 = 20

Present value = $130,000 *PV factor at 5% for 20th period

= $130,000 * 0.37689 = $48,996

Solution d:

Present value = $260,000 * PV factor at 10% for 10th period

= $260,000 * 0.38554 = $100,240